# Multiscale contact mechanics for elastoplastic contacts

### v1

@@ 5  Discussion
A key result of our study is that we demonstrate how the boundary conditions $P(\sigma,\zeta)\to 0$ as $\sigma\to 0^{+}$ and $\sigma\to\sigma_{\rm P}^{-}$, which are assumed in the analytical theory, naturally emerge from the boundary element simulations. These conditions are not imposed explicitly; rather, they follow from the way plasticity is implemented. Specifically, nodes where $\sigma>\sigma_{\rm P}$ are classified as plastically loaded and remain in contact without bearing additional load, while nodes with $0<\sigma\leq\sigma_{\rm P}$ and $\sigma=0$ are considered elastically loaded and non-contacting, respectively. While this setup constrains the stress distribution to lie within the physically admissible range, it does not, a priori, enforce that $P(\sigma,\zeta)$ vanishes continuously at the boundaries. Nevertheless, the numerical results shown in Fig. 5 clearly exhibit this behaviour, thereby validating the theoretical boundary conditions in the continuum limit. This contrasts with the work of Xu et al. [4], who assumed these boundary conditions without providing a rigorous justification in the elastoplastic case.
An important parameter in the elastoplastic theory is the penetration hardness, and here we will discuss it in some detail. Indentation hardness can be measured using indenters of various sizes and shapes. Here, we define the penetration hardness as the applied force divided by the projected indentation area. In general, the penetration hardness depends on the shape of the indenter (because of the different form of the stress field and plastic deformation)[18]. Macroscopic tests usually use spherical or pyramidal shapes. Most rough surfaces have “smooth” asperities like spherical cups (sharp spike-like structures are unlikely as they would be removed during the surface preparation, such as polishing), and we believe using a spherical indenter (as in the Brinell hardness test) gives a penetration hardness most suitable for contact mechanics calculations.
@@
There are several other possible origins of why the penetration hardness may depend on the length scale. Many solids have modified surface layers (e.g. oxide films on metals), and the indentation hardness may be very different if the size of the indentation is smaller or larger than the layer thickness.
Figure 6: (a) The surface height profile $h(x)$ after squeezing a steel ball (radius $R=2\ {\rm cm}$) with smooth surface against a sandblasted aluminum surface. The middle line shows the height profile from inside the indentation, and the bottom line shows the height profile after removing the macroscopic curvature. (b) The red lines are two height profiles $h(x)$ of the aluminum surface before indenting it with the steel ball, and the blue line is the height profile from inside the indentation [bottom linescan in (a)]. Adapted from Ref.[13].
Even without strain hardening or modified surface layers, the effective penetration hardness may depend on the amount of plastic deformation. This is illustrated in Fig. 6, which shows the indentation of a nominal flat surface of a solid (here sandblasted aluminum) with surface roughness by a rigid and smooth spherical indenter[13]. In the most naive picture of plastic deformation, one would expect all the asperities within the indented (spherical cup) region to be plastically flattened. This result follows from the fact that the local pressure acting on an asperity, if not flattened, would be larger than the indentation hardness as given by the normal force divided by the projected (macroscopic) indentation area. However, it has been known for a long time (and illustrated in Fig. 6) that the asperities do not flatten out perfectly, but only the upper part of some fraction of the asperities flatten in such a way that the flattened area typically occupies the macroscopic indentation area[27, 28, 29].
This persistence of asperities in indentation experiments can be interpreted as resulting from a indentation hardness, which increases as the length scale (or indentation size) decreases, but a more likely explanation is related to plasticity mechanics of asperity interaction as proposed in [27, 28, 30, 31, 32]. That is, qualitatively, as an asperity becomes strongly plastically deformed, the stress field approaches a hydrostatic stress, and the asperity therefore becomes resistant to further plastic deformation. This is not a length scale effect and cannot be simply included as a dependency of the penetration hardness on the length scale of the roughness or magnification.
@@
6  Summary and conclusion
We have tested Persson’s multiscale contact mechanics theory for elastoplastic solids, assuming a constant penetration hardness. Using a spectral boundary element method, we simulated contact between rigid, randomly rough surfaces and an elastic–perfectly plastic half-space with a flat surface. The theory’s predictions for elastic, plastic, and total contact areas are in excellent quantitative agreement with the numerical results. Moreover, the simulations confirm that the boundary conditions assumed in the theory, i.e., that the stress probability distribution vanishes at both zero and yield stress, arise naturally in the numerical model. These findings validate the theoretical framework for systems with constant hardness and support its use in practical contact mechanics applications. While scale-dependent hardness may be important in some cases, the constant-hardness model captures the dominant behavior in a wide range of scenarios.

### v3

@@ V Discussion
A key result of our study is that we demonstrate how the boundary conditions $P(\sigma,\zeta)\to 0$ as $\sigma\to 0^{+}$ and $\sigma\to\sigma_{\rm P}^{-}$, which are assumed in the analytical theory, naturally emerge from the boundary element simulations. These conditions are not imposed explicitly; rather, they follow from the way plasticity is implemented. Specifically, nodes where $\sigma>\sigma_{\rm P}$ are classified as plastically loaded and remain in contact without bearing additional load, while nodes with $0<\sigma<\sigma_{\rm P}$ and $\sigma=0$ are considered elastically loaded and non-contacting, respectively. While this setup constrains the stress distribution to lie within the physically admissible range, it does not, a priori, enforce that $P(\sigma,\zeta)$ vanishes continuously at the boundaries. Nevertheless, the numerical results shown in Fig. 5 clearly exhibit this behavior, thereby validating the theoretical boundary conditions in the continuum limit.
Xu et al. [5] derived a simple closed-form expression for the stress probability distribution, $P(\sigma,\zeta)$, that satisfies these boundary conditions. They also presented a consistency argument in support of them. However, because their closed-form solution was itself constructed by imposing these conditions, the resulting argument does not constitute an independent derivation of the boundary conditions $P(\sigma,\zeta)\to 0$ as $\sigma\to 0^{+}$ and $\sigma\to\sigma_{\rm P}^{-}$ in Persson’s elastoplastic contact mechanics theory. Our BEM results demonstrate these limits naturally without such an assumption.
An important parameter in the elastoplastic theory is the penetration hardness and here will discuss it in some detail. Indentation hardness can be measured using indenters of various sizes and shapes. Here we define the penetration hardness as the applied force divided by the projected indentation area. In general, the penetration hardness depends on the shape of the indenter (because of the different form of the stress field and plastic deformation) [24]. Macroscopic tests usually use spherical or pyramidal shapes. Most rough surfaces, when observed at any given magnification, have “smooth” asperities like spherical cups (sharp spike-like structures are unlikely as they would be removed during the surface preparation, such as polishing), and we believe using a spherical indenter (as in the Brinell hardness test) gives a penetration hardness most suitable for contact mechanics calculations.
@@
Even without strain hardening the effective penetration hardness may depend on the amount of plastic deformation. Thus, it has been known for a long time that the asperities do not flatten out perfectly, but only the upper part of some fraction of the asperities flatten in such a way that the flattened area typically occupies half of the macroscopic indentation area [33, 34, 35]. This persistence of asperities in indentation experiments can be interpreted as resulting from an indentation hardness, which increases as the length scale (or indentation size) decreases, but a more likely explanation is related to plasticity mechanics of asperity interaction as proposed in [33, 34, 36, 37, 38]. That is, qualitatively, as an asperity becomes strongly plastically deformed, the stress field approaches a hydrostatic stress, and the asperity therefore becomes resistant to further plastic deformation. This is not a length scale effect and cannot be simply included as a dependency of the penetration hardness on the length scale of the roughness or magnification.
@@
VI Summary and conclusion
We have tested Persson’s multiscale contact mechanics theory for elastoplastic solids, assuming a constant penetration hardness. Using a spectral boundary element method, we simulated contact between a elastic–perfectly plastic half-space with randomly rough surfaces, and a rigid substrate with a flat surface. The theory’s predictions for elastic, plastic, and total contact areas are in excellent quantitative agreement with the numerical results. Moreover, the simulations confirm that the boundary conditions assumed in the theory, i.e., that the stress probability distribution vanishes at both zero and yield stress, arise naturally in the numerical model. These findings validate the theoretical framework for systems with constant hardness and support its use in practical contact mechanics applications. While scale-dependent hardness may be important in some cases, the constant-hardness model captures the dominant behavior in a wide range of scenarios.